what is the slope of a line that goes through the points (2,5) and (8,6)
Answer by checkley77(12844)

(2,5) and (8,6)
Slope=(y2-y1)/(x2-x1)\
Slope=(6-5)/(8-2)
Slope=1/6 ans. for the slope.